OpenStudy (angelax0):

Frank drove to the mountains last weekend. There was heavy traffic on the way there, and the trip took 7 hours. When Frank drove home, there was no traffic and the trip only took 5 hours. If his average rate was 18 miles per hour faster on the trip home, how far away does Frank live from the mountains? Do not do any rounding.

OpenStudy (wolf1728):

rate = distance/time let distance=x rate +18 = x/5 rate = x/7 a) x = 5 * rate + 90 b) x = 7*rate multiplying a) by -1 a) -x = -5rate -90 b) x = 7rate Adding the equations 2 rate = 90 rate = 45 rate + 18 = 63 distance = 7*45 = 315

OpenStudy (wolf1728):

angelax0 - what do you think?

OpenStudy (angelax0):

thank you

OpenStudy (wolf1728):

u r welcome